# Modular Jump Gaussian Processes

### v2

@@ 1 Introduction
The top left panel of Figure 1 shows an ordinary/stationary GP fit, as reviewed in Section 2.1. Three issues are of note here: the GP is mistaking signal for noise, overestimating variance, and smoothing over the jumps. It seems to only have detected signal from the jumps, not from the sinusoid. Uncertainty is uniform because inputs are uniform and the GP assumes stationarity. Now turn to the top right panel, showing a local (approximate) GP (LAGP), reviewed in Section 2.2. For now we remark that a LAGP is basically an ordinary GP where training data far from the testing location(s) is not used. Less data means faster calculations and increased “reactivity” if dynamics change in the input space. Notice how LAGP shows an improvement over the ordinary GP (top left) for $x<40$ and $x>80$. Signal is detected there, and variance is much reduced. But this comes at the expense of accuracy in the middle region, and near the “jumps” on either side. Uncertainty here is inflated relative to the stationary GP fit. Importantly, LAGP uses a fixed-sized neighborhood of locality everywhere in the input space which we shall explain in more detail in Section 2.2. Our proposed approach improves upon this by dynamically selecting the optimal neighborhood size throughout the input space as described in Section 3.
@@
The lower panels of Figure 1 highlight our main contribution. Focus first on the lower-left. Suppose we were to cluster the observations marginally. We ignore $(x_{1},\dots,x_{N})$ and use an unsupervised learning scheme to assign (one of two) labels to each of $(y_{1},\dots,y_{N})$. In this 1d example, most clustering methods will easily partition these $y$-values into components around 0 and 10. We use model-based clustering via EM for this, as described in Section 4.2, but suggest several variations and other approaches that often work just as well. Cluster memberships in hand, e.g., $c_{i}=1$ for $y_{i}$ around 0 and $c_{j}=2$ for $y_{j}$ around 10, suppose we were to learn a classification rule spatially for those pairs $(x_{1},c_{1}),\dots,(x_{N},c_{N})$, i.e., bringing the $x$’s back in with supervised learning. There are again many choices for this; we prefer a classification GP (CGP). This allows us to predict a probability of class 2 at $x$, in-sample at ($x_{1},\dots,x_{N}$) yielding $(p_{1},\dots,p_{N})$, or out-of sample for testing as $p(x)$. We then treat the class predictions $(p_{1}\dots p_{N})$, derived from the training data, as an input feature in our GP fit. That is, take the same GP from the top-left panel, but train it on input tuple $(x_{i},p_{i})$ and output $y_{i}$, for $i=1,\dots N$. When predicting, augment any testing $x$ with $p(x)$ for tuple $(x,p(x))$. This global GP with the new feature is visualized in the lower left panel of Figure 1. Notice how the fit tracks the jumps better (compared to the top left), and it is more confident everywhere, but it still struggles to pick up on the sinusoid. The idea of learning an auxiliary input feature to accommodate nonstationarity is not new (Bornn et al.,, 2012) but, like Park et al., (2022) for JGP, Bornn et al., perform joint inference for the GP and the latent quantity, which can be cumbersome. A key feature of our approach is that we do everything in a pre-processing step with off-the-shelf libraries. This both increases the flexibility of our approach and makes it more widely accessible.
@@ 2.1 Gaussian Processes
As a shorthand for later, let $\mathrm{GP}\left(\mathcal{X};X_{N},Y_{N}\right)$ indicate the use of Eqs. (24) to represent predictions at $\mathcal{X}$ via MLE hyperparameters using training data $X_{N}$ and $Y_{N}$. We refer to this as a “global GP” since all training data are used. Such a fit is valid for prediction at any $\mathcal{X}$ in the input space. This is in contrast to a GP fit using a subset of the training data, which we shall review next. The GP shown in the top left panel of Figure 1 is a global GP, where the solid yellow line is $\mu_{N}(\mathcal{X})$ and the dashed yellow line is a 95% prediction interval (PI) calculated using the diagonal entries of $\Sigma_{N}(\mathcal{X})$.
@@ 2.2 Local Approximate Gaussian Processes
The core idea behind LAGP is that fitting a GP to $X_{n}(x)\subset X_{N}$ with $n\ll N$, along with corresponding $Y_{n}$, is both faster, and more reactive to local dynamics around $x$. We encapsulate that in shorthand as $\text{LAGP}_{n}(x)=\text{GP}\left(x;X_{n}(x),Y_{n}\right)$ after optimizing any hyperparameters in $\Sigma(X_{n}(x))$, and $\text{LAGP}_{n}(\mathcal{X})$ for $m$ separate applications covering all of $\mathcal{X}$. The top right panel of Figure 1 shows $\text{LAGP}_{25}(\mathcal{X})$, where the red solid line denotes $\mu_{N}(\mathcal{X})$ and the red dashed line denotes a 95% PI. Since calculations are computationally exclusive for each ${x}\in\mathcal{X}$, we don’t get a full covariance structure $\Sigma_{N}(\mathcal{X})$, only a diagonal one with entries $\sigma^{2}_{x}$. The upside, though, is that the calculations are massively parallelizable (Gramacy et al.,, 2014).
@@
Figure 2: Three possible neighborhood choices for $x=(0.04,0.023)$ (the blue dot). The two colors (red and yellow) are two levels of the response, with a manifold of discontinuity at $\sin(x_{1})=x_{2}$. The orange and yellow dots are $X_{N}$ and the black dots are three different choices for $X_{n}(x)$.
@@ 3 Optimal Neighborhood-size LAGP
The image on the left is formed by evaluating the test function on a dense $101\times 101$ grid in $0.5,0.5]^{2}$. All responses ($y$-values) are deterministic. Observe that there is smooth color variation within the (mostly) red and yellow regions, revealing a strong signal locally (light reds and dark yellows, respectively), even away from the “jumps”. The histogram on the right helps understand the magnitudes of these signals. Notice that red and yellow regimes are have comparable within-color variability and they are well-separated, with no values between 0.4 and 0.6.
@@ 3.1 Searching via validation
Consider a single predictive location $x$ and a set of possible neighborhood sizes $n\equiv n(x)\in\{n_{\min},\dots,n_{\max}\}$. We wish to estimate an optimal $\hat{n}(x)$ by some criterion, where $X_{n_{\min}}(x)\subseteq X_{\hat{n}}(x)\subseteq X_{n_{\max}}(x)\subseteq X_{N}$, as solved by LAGP [Section 2.2]. Let $\mu_{n}(x)$ denote the predictive mean under some $X_{n}(x)$, i.e., following Eq. (4), but conditional on $(X_{n}(x),Y_{n})$. This prediction has some accuracy, say measured by squared error $\mathrm{SE}(y,\mu_{n}(x))$ where $y$ is the true, but unknown value of the response at $x$. Generically, $\mathrm{SE}(y,\hat{y})=(y-\hat{y})^{2}$. One of the main contributions of Gramacy and Apley, (2015) involves arguing that NN-constructed $X_{n}(x)$, conditional on a neighborhood size $n$, approximately minimizes the the expectation of this quantity (i.e., the mean-squared error) for $y$ following a stationary GP.
Here, we wish to extend that to choosing the best $\hat{n}$. However, we can’t use MSE under a stationary GP for $y$. There are two reasons for this. One is practical: this is not the case we’re interested in – our jump processes are nonstationary. The other is more fundamental. If the process is stationary, then MSE is a decreasing function of the the subset size $n$. In other words, you want to condition on as much data as possible. Since large $n$ is generally intractable, Gramacy and Apley, recommended choosing $n$ as large as computational constraints allow. However, they observed that performance varied with $n$ when the data exhibited telltale signs of nonstationarity, leaving an open question about what to do in practice. So instead of averaging SE, we create a validation exercise which allows us to plug-in an actual $y_{i}$ value from the training data as a proxy.
Let $x_{*}$ be the input in $X_{N}$ that is closest (in terms of Euclidean distance) to $x$. Note that this is also the closest point to $x$ in any $X_{n}(x)$ for $n\geq 1$ and that $x_{*}\equiv X_{1}(x)$. We prefer the simpler $x_{*}$, which is easier on the eyes than $X_{1}(x)$. Analogously, let $y_{*}$ be the element in $Y_{N}$ corresponding to $x_{*}$. We use $y_{*}$ as a proxy for the value of $y(x)$, the mean response at new location $x$, which is unknown. Next, let $X_{n}(x_{*})$ denote an $n$-sized neighborhood for predicting at $x_{*}$, i.e., the $n$-nearest points to $x_{*}$ in $X_{N}$, excluding $x_{*}$. Using these quantities, and for any neighborhood size $n$, one may fit $\text{LAGP}_{n}\left(x_{*};X_{n}(x_{*}),Y_{n}\right)$, furnishing a prediction for $y_{*}$. Specifically, let $\mu_{n}(x_{*})$ denote the predictive mean of $x_{*}$ calculated with neighborhood size $n$. Suppose we do this for each $n$ of interest, and take the best one by squared error:
@@
These updates, and the computational cost analysis, presume that $\Sigma_{n-1}$ and $\Sigma_{n}$ utilize the the same hyperparameters, like lengthscales $\theta_{k}$, which may not be ideal. In particular, shorter lengthscales will generally be preferred for smaller $n_{\min}$-sized neighborhoods whose squared distances (3) are short. Whereas when $n$ is closer to $n_{\max}$, including longer squared distances, longer lengthscales may be preferred. However, an MLE calculation (2) after each update, as new $x_{n}$ arrives, would incur cubic costs and result in a quartic overall computational complexity, which isn’t ideal either. Recognizing that an additional data point is of greater value to the likelihood, say via its Fisher information, when $n$ is small as opposed to large, we deploy a logarithmic MLE-updating schedule so that cubic MLE costs are incurred much less often as $n$ increases. Any sub-linear updating schedule will keep computation cubic.
@@ 3.2 An Illustration
Now consider how all 10K-odd neighborhoods, over the whole grid $\mathcal{X}$, compare to the laGP software default of $n=50$. Figure 5 shows how that choice (between the vertical lines) is not advantageous for the Phantom data. Many smaller, but also many much larger larger neighborhoods are estimated by OLAGP, and with much greater frequency. The figure also provides neighborhoods for the new method (“MJGP”) discussed next in Section 4.
Before turning to that, we note anecdotally that OLAGP performs worst at or near the manifold of discontinuity (See Appendix A.1). This is because it cannot choose $\hat{n}(x)$ small enough to exclude points from the opposite side of the manifold without maintaining a viable minimum ($n_{\min}=12$) for a GP fit. Observe how the neighborhood for $(0.39,0.34)$ contains several points in the red region. At $\hat{n}(x)=14$ for that location, the neighborhood can’t shrink much further. Consequently, the resulting mean predictions (4) are weighted averages of red- and yellow-region response values. This will yield predictions between 0.4 and 0.6, which we know from Figure 3 are inaccurate. Shrinking neighborhoods farther, even below $n_{\min}=12$, is a double-edged sword: less data means greater reactivity, but also lower information content. It would be better to allow a more flexible notion of neighborhood, beyond simple Euclidean distance in $X_{N}$. Next, we suggest how that can be done without fundamentally altering the LAGP setup by introducing a latent variable. To foreshadow, observe that the “MJGP” neighborhood sizes summarized in Figure 5 show fewer small neighborhoods, and fewer large neighborhoods compared to OLAGP.
@@ 4.1 Identifying and predicting levels
Data $Y_{N}$ from a jump process can often be partitioned marginally (i.e., $y$ without $x$), which can inform a more complex spatial partition over $x$. To see this, consider again the histogram of the marginal $Y_{N}$ for the Phantom data [Figure 3, right]. There are two, roughly symmetric, well-separated densities, centered at 0.2 and 0.8. These modes are so well separated you can eyeball it, say at $y=0.5$.
@@
Let $c_{i}\in\{1,2\}$, $i=1,\dots,N$ denote the level learned for each $y_{i}\in Y_{N}$ via model based clustering, with vector $C_{N}$ holding all $N$ levels. For example, we could have $c_{i}=2$ correspond to “yellow” and 1 for “red”, or vice-versa. Then, pair these with inputs $X_{N}$ to form $(X_{N},C_{N})$, comprising “data” for a spatial classification task. Here we shall write $\hat{f}(\cdot)=\mathrm{class}(X_{N},C_{N})$, generically, for the predictive equations of a fitted, spatial classification. One option is (basis-expanded) linear logistic regression, or logistic GP classification (e.g Rasmussen and Williams,, 2006, Chapter 3). Any implementation works, although the accuracy of a particular method on a particular data set should be tested before proceeding; our preferred implementation is detailed in Section 5. We may then use $\hat{f}$ to predict component membership at testing locations $\mathcal{X}$: $\hat{f}(x)=\mathbb{P}(y(x)=1)$ for any $x\in\mathcal{X}$. Predicted probabilities for the Phantom data are shown in the left panel of Figure 7.

### v3

@@ 1 Introduction
The top left panel of Figure 1 shows an ordinary/stationary GP fit, as reviewed in Section 2.1. Three issues are of note here: the GP is mistaking signal for noise, overestimating variance, and smoothing over the jumps. It seems to only have detected signal from the jumps, not from the sinusoid. Uncertainty is uniform because inputs are uniform and the GP assumes stationarity. Now turn to the top right panel, showing a local (approximate) GP (LAGP), reviewed in Section 2.2. For now we remark that a LAGP is basically an ordinary GP where training data far from the testing location(s) is not used. Less data means faster calculations and increased “reactivity” if dynamics change in the input space. Notice how LAGP shows an improvement over the ordinary GP (top left) for $x<40$ and $x>80$. Signal is detected there, and variance is much reduced. But this comes at the expense of accuracy in the middle region, and near the “jumps” on either side. Uncertainty here is inflated relative to the stationary GP fit. Importantly, LAGP uses a fixed-sized neighborhood of locality ($n=25$) everywhere in the input space which we shall explain in more detail in Section 2.2. Our proposed approach improves upon this by dynamically selecting the optimal neighborhood size throughout the input space as described in Section 3.
@@
The lower panels of Figure 1 highlight our main contribution. Focus first on the lower-left. Suppose we were to cluster the observations marginally. We ignore $(x_{1},\dots,x_{N})$ and use an unsupervised learning scheme to assign (one of two) labels to each of $(y_{1},\dots,y_{N})$. In this 1d example, most clustering methods will easily partition these $y$-values into components around 0 and 10. We use model-based clustering via EM for this, as described in Section 4.2, but suggest several variations and other approaches that often work just as well. Using cluster memberships, e.g., $c_{i}=1$ for $y_{i}$ around 0 and $c_{j}=2$ for $y_{j}$ around 10, suppose we were to learn a classification rule spatially for pairs $(x_{1},c_{1}),\dots,(x_{N},c_{N})$, i.e., reincorporating the inputs $(x_{1},\dots,x_{N})$ with supervised learning. There are again many choices for this; we prefer a classification GP (CGP). This allows us to predict a probability of class 2 at $x$, in-sample at ($x_{1},\dots,x_{N}$) yielding $(p_{1},\dots,p_{N})$, or out-of sample for testing as $p(x)$. We then treat the class predictions $(p_{1}\dots p_{N})$, derived from the training data, as an input feature in our GP fit. That is, take the same GP from the top-left panel, but train it on input tuple $(x_{i},p_{i})$ and output $y_{i}$, for $i=1,\dots N$. When predicting, augment any testing $x$ with $p(x)$ for tuple $(x,p(x))$. This global GP with the new feature is visualized in the lower left panel of Figure 1. Notice how the fit tracks the jumps better (compared to the top left), and it is more confident everywhere, but it still struggles to pick up on the sinusoid. The idea of learning an auxiliary input feature to accommodate nonstationarity is not new (Bornn et al.,, 2012) but, like Park et al., (2022) for JGP, Bornn et al., perform joint inference for the GP and the latent quantity, which can be cumbersome. A key feature of our approach is that we do everything in a pre-processing step with off-the-shelf libraries. This both increases the flexibility of our approach and makes it more widely accessible.
@@ 2.1 Gaussian Processes
As a shorthand for later, let $\mathrm{GP}\left(\mathcal{X};X_{N},Y_{N}\right)$ indicate the use of Eqs. (24) to represent predictions at $\mathcal{X}$ via MLE hyperparameters using training data $X_{N}$ and $Y_{N}$. We refer to this as a “global GP” since all training data are used. Such a fit is valid for prediction at any $\mathcal{X}$ in the input space. This is in contrast to a GP fit using a subset of the training data, which we shall review next. The GP shown in the top left panel of Figure 1 is a global GP, where the solid purple line is $\mu_{N}(\mathcal{X})$ and the dashed purple line is a 95% prediction interval (PI) calculated using the diagonal entries of $\Sigma_{N}(\mathcal{X})$.
@@ 2.2 Local Approximate Gaussian Processes
The core idea behind LAGP is that fitting a GP to $X_{n}(x)\subset X_{N}$ with $n\ll N$, along with corresponding $Y_{n}$, is both faster, and more reactive to local dynamics around $x$. We encapsulate that in shorthand as $\text{LAGP}_{n}(x)=\text{GP}\left(x;X_{n}(x),Y_{n}\right)$ after optimizing any hyperparameters in $\Sigma(X_{n}(x))$, and $\text{LAGP}_{n}(\mathcal{X})$ for $m$ separate applications covering all of $\mathcal{X}$. The top right panel of Figure 1 shows $\text{LAGP}_{25}(\mathcal{X})$, where the red solid line denotes $\mu_{N}(\mathcal{X})$ and the red dashed line denotes a 95% PI. Since calculations are computationally exclusive for each ${x}\in\mathcal{X}$, we do not get a full covariance structure $\Sigma_{N}(\mathcal{X})$, only a diagonal one with entries $\sigma^{2}_{x}$. The upside, though, is that the calculations are massively parallelizable (Gramacy et al.,, 2014).
@@
Figure 2: Three possible neighborhood choices for $x=(0.04,0.023)$ (the blue dot). The two colors (orange and yellow) are two levels of the response, with a manifold of discontinuity at $\sin(x_{1})=x_{2}$. The orange and yellow dots are $X_{N}$ and the black dots are three different choices for $X_{n}(x)$.
@@ 3 Optimal Neighborhood-size LAGP
The image on the left is formed by evaluating the test function on a dense $101\times 101$ grid in $0.5,0.5]^{2}$. All responses ($y$-values) are deterministic. Observe that there is smooth color variation within the (mostly) red and yellow regions, revealing a strong signal locally (light reds and dark yellows, respectively), even away from the “jumps”. The histogram on the right helps understand the magnitudes of these signals. Notice that red and yellow regimes have comparable within-color variability and they are well-separated, with no values between 0.4 and 0.6.
@@ 3.1 Searching via validation
Consider a single predictive location $x$ and a set of possible neighborhood sizes $n\equiv n(x)\in\{n_{\min},\dots,n_{\max}\}$. We wish to estimate an optimal $\hat{n}(x)$ by some criterion, where $X_{n_{\min}}(x)\subseteq X_{\hat{n}}(x)\subseteq X_{n_{\max}}(x)\subseteq X_{N}$, as solved by LAGP [Section 2.2]. Let $\mu_{n}(x)$ denote the predictive mean under some $X_{n}(x)$, i.e., following Eq. (4), but conditional on $(X_{n}(x),Y_{n})$. This prediction has some accuracy, say measured by squared error $\mathrm{SE}(y,\mu_{n}(x))$, where $y$ is the true but unknown value of the response at $x$. Generically, $\mathrm{SE}(y,\hat{y})=(y-\hat{y})^{2}$. One of the main contributions of Gramacy and Apley, (2015) involves arguing that NN-constructed $X_{n}(x)$, conditional on a neighborhood size $n$, approximately minimizes the expectation of this quantity (i.e., the mean-squared error) for $y$ following a stationary GP.
Here, we wish to extend that to choosing the best $\hat{n}$. However, we cannot use MSE under a stationary GP for $y$. There are two reasons for this. One is practical: this is not the case we are interested in – our jump processes are nonstationary. The other is more fundamental. If the process is stationary, then MSE is a decreasing function of the subset size $n$. In other words, you want to condition on as much data as possible. Since large $n$ is generally intractable, Gramacy and Apley, recommended choosing $n$ as large as computational constraints allow. However, they observed that performance varied with $n$ when the data exhibited telltale signs of nonstationarity, leaving an open question about what to do in practice. So instead of averaging SE, we create a validation exercise which allows us to plug-in an actual $y_{i}$ value from the training data as a proxy.
Let $x_{*}$ be the input in $X_{N}$ that is closest (in terms of Euclidean distance) to $x$. Note that this is also the closest point to $x$ in any $X_{n}(x)$ for $n\geq 1$ and that $x_{*}\equiv X_{1}(x)$. For ease of notation, we shall henceforth refer to $X_{1}(x)$ as $x_{*}$. Analogously, let $y_{*}$ be the element in $Y_{N}$ corresponding to $x_{*}$. We use $y_{*}$ as a proxy for the value of $y(x)$, the mean response at new location $x$, which is unknown. Next, let $X_{n}(x_{*})$ denote an $n$-sized neighborhood for predicting at $x_{*}$, i.e., the $n$-nearest points to $x_{*}$ in $X_{N}$, excluding $x_{*}$. Using these quantities, and for any neighborhood size $n$, one may fit $\text{LAGP}_{n}\left(x_{*};X_{n}(x_{*}),Y_{n}\right)$, furnishing a prediction for $y_{*}$. Specifically, let $\mu_{n}(x_{*})$ denote the predictive mean of $x_{*}$ calculated with neighborhood size $n$. Suppose we do this for each $n$ of interest, and take the best one by squared error:
@@
These updates, and the computational cost analysis, presume that $\Sigma_{n-1}$ and $\Sigma_{n}$ utilize the same hyperparameters, like lengthscales $\theta_{k}$, which may not be ideal. In particular, shorter lengthscales will generally be preferred for smaller $n_{\min}$-sized neighborhoods whose squared distances (3) are short. Whereas when $n$ is closer to $n_{\max}$, including longer squared distances, longer lengthscales may be preferred. However, an MLE calculation (2) after each update, as new $x_{n}$ arrives, would incur cubic costs and result in a quartic overall computational complexity, which is not ideal either. Recognizing that an additional data point is of greater value to the likelihood, say via its Fisher information, when $n$ is small as opposed to large, we deploy a logarithmic MLE-updating schedule so that cubic MLE costs are incurred much less often as $n$ increases. Any sub-linear updating schedule will keep computation cubic.
@@ 3.2 An Illustration
Now consider how all 10,201 neighborhoods, over the whole grid $\mathcal{X}$, compare to the laGP software default of $n=50$. Figure 5 shows how that choice (between the vertical lines) is not advantageous for the Phantom data. Many smaller, but also many much larger neighborhoods are estimated by OLAGP, and with much greater frequency. The figure also provides neighborhoods for the new method (“MJGP”) discussed next in Section 4.
Before turning to that, we note anecdotally that OLAGP performs worst at or near the manifold of discontinuity (See Appendix B.1). This is because it cannot choose $\hat{n}(x)$ small enough to exclude points from the opposite side of the manifold without maintaining a viable minimum ($n_{\min}=12$) for a GP fit. Observe how the neighborhood for $(0.39,0.34)$ contains several points in the red region. At $\hat{n}(x)=14$ for that location, the neighborhood cannot shrink much further. Consequently, the resulting mean predictions (4) are weighted averages of red- and yellow-region response values. This will yield predictions between 0.4 and 0.6, which we know from Figure 3 are inaccurate. Shrinking neighborhoods farther, even below $n_{\min}=12$, is a double-edged sword: less data means greater reactivity, but also lower information content. It would be better to allow a more flexible notion of neighborhood, beyond simple Euclidean distance in $X_{N}$. Next, we suggest how that can be done without fundamentally altering the LAGP setup by introducing a latent variable. To foreshadow, observe that the “MJGP” neighborhood sizes summarized in Figure 5 show fewer small neighborhoods, and fewer large neighborhoods compared to OLAGP.
@@ 4.1 Identifying and predicting levels
Data $Y_{N}$ from a jump process can often be partitioned marginally (i.e., $y$ without $x$), which can inform a more complex spatial partition over $x$. To see this, consider again the histogram of the marginal $Y_{N}$ for the Phantom data [Figure 3, right]. There are two, roughly symmetric, well-separated densities, centered at 0.2 and 0.8. These modes are so well separated you can partition it visually, say at $y=0.5$.
@@
Let $c_{i}\in\{1,2\}$, $i=1,\dots,N$ denote the level learned for each $y_{i}\in Y_{N}$ via model based clustering, with vector $C_{N}$ holding all $N$ levels. For example, we could have $c_{i}=2$ correspond to “yellow” and 1 for “red”, or vice-versa. Then, pair these with inputs $X_{N}$ to form $(X_{N},C_{N})$, comprising “data” for a spatial classification task. Here we shall write $\hat{f}(\cdot)=\mathrm{class}(X_{N},C_{N})$, generically, for the predictive equations of a fitted, spatial classification. One option is (basis-expanded) linear logistic regression, or logistic GP classification (e.g Rasmussen and Williams,, 2006, Chapter 3). Any classification method works, although the accuracy of a particular method on a particular data set should be tested, for example by evaluating classification accuracy on $X_{N}$, before proceeding. Our preferred implementation is detailed in Section 5. We may then use $\hat{f}$ to predict component membership at testing locations $\mathcal{X}$: $\hat{f}(x)=\mathbb{P}(y(x)=1)$ for any $x\in\mathcal{X}$. Predicted probabilities for the Phantom data are shown in the left panel of Figure 7 with the manifold of discontinuity overlaid.